# ColNet: Collaborative Optimization in Decentralized Federated Multi-task Learning Systems

### v1

@@ 3.4 Learning Phase
6:     for $k\in K$ do $\triangleright$ Intra-task Aggregation
7:         Send $\mathbf{w}^{B}$ to $k$
8:         Receive $\mathbf{w}^{B}_{k}$ from $k$
9:     end for
10:     $\mathbf{w}^{B}\leftarrow\text{Avg}\bigl{(}\{\mathbf{w}^{B}_{k}\}_{k\in K}\cup\{\mathbf{w}^{B}\}\bigr{)}$triangleright Eq. (2)
11:     if $isLeader$ then
12:         $\Delta\bar{\mathbf{w}}^{B}\leftarrow\mathbf{w}^{B}-\mathbf{w}^{B-1}$
13:         for $l\in L$ do $\triangleright$ Cross-task Aggregation
14:              Send $\Delta\bar{\mathbf{w}}^{B}$ to $l$
15:              Receive $\Delta\bar{\mathbf{w}}^{B}_{l}$ from $l$
16:         end for
17:         $\tilde{U}\leftarrow\text{HCA}\!\bigl{(}\Delta\bar{\mathbf{w}}^{B},\{\Delta\bar{\mathbf{w}}^{B}_{l}\}\bigr{)}$triangleright Eq. (4)
18:         $\mathbf{w}^{B}\leftarrow\mathbf{w}^{B}+\tilde{U}$
19:         for $k\in K$ do
20:              Send $\mathbf{w}^{B}$ to $k$
21:         end for
22:     else
23:         $\mathbf{w}^{B}\leftarrow\text{Wait for leader update}$
24:     end if
25:end for
Intra-Task Aggregation.
Within a task group, after each local training round, every node exchanges its updated backbone parameters with all other nodes in the same task group, then averages the collected backbones to update its own. Formally, if there are $K$ nodes in a group and $\mathbf{w}_{i}^{B}$ denotes the backbone parameters of node $i$, then the intra-task aggregation can be expressed as:
$\bar{\mathbf{w}}^{B}=\frac{1}{K}\sum_{i=1}^{K}\mathbf{w}^{B}_{i}.$ (2)
Cross-Task Aggregation.
Cross-task aggregation is more challenging due to potential conflicts among updates originating from distinct tasks. These conflicts arise when backbone parameters diverge significantly, leading to dominant or opposing gradients that degrade training. Inspired by FedHCA2 Lu et al. (2024), the cross-task aggregation in ColNet adopts a modified version of their gradient-alignment approach in decentralized learning environments.
ColNet uses a gradient term that maximizes the inner product between conflicting updates, effectively reducing gradient conflict. The core idea is to find an aggregated update $\tilde{U}$ (for the group leaders’ backbone) that remains close to the average update $\Delta\bar{\mathbf{w}}^{B}=\tfrac{1}{N}\sum_{i=1}^{N}\Delta\mathbf{w}i^{B}$ while minimizing inter-task conflicts:
@@
where $c\in[0,1)$ regulates how far $\tilde{U}$ can deviate from $\Delta\bar{\mathbf{w}}^{B}$.
By introducing a Lagrangian formulation and simplifying, the final update is shown to be:
$\tilde{U}\;=\;\Delta\bar{\mathbf{w}}^{B}\;+\;\frac{\sqrt{\phi}}{\|U_{w}\|}\,U_{w},$ (4)
where $U_{w}=\tfrac{1}{N}\sum_{i=1}^{N}w_{i}\Delta\mathbf{w}^{B}_{i}$ and $\phi=c^{2}\|\Delta\bar{\mathbf{w}}^{B}\|^{2}$. This formulation balances each leader’s update direction with the average update, mitigating conflicts while maintaining proximity to the mean. Upon computing $\tilde{U}$, each group leader updates its backbone and shares it with all group members, thereby propagating the conflict-reduced update throughout its task group.
4 Deployment and Evaluation
This section details the deployment and experimental evaluation of ColNet, covering the targeted multi-task learning scenarios, the chosen evaluation metrics, and the resulting performance outcomes.
4.1 Scenarios, Datasets, and Models
@@
Label heterogeneity arises when clients train on different subsets of class labels, which requires splitting the model into a shared backbone and task-specific layers to protect privacy and maintain performance. Clients with the same labels form a single task group, sharing only the backbone parameters. The CIFAR-10 dataset Krizhevsky et al. (2009) was selected to evaluate label heterogeneity, where each image belongs to one of ten classes. Two task groups were formed: animals (bird, cat, deer, dog, frog, horse) and objects (airplane, automobile, ship, truck) classification. Each group contains three clients, and only the backbone parameters are shared within each group.
@@
All clients train a ResNet-18 He et al. (2016) backbone and separate task layers to accommodate different output dimensions (four classes for objects, six for animals). A Cross-Entropy loss and SGD optimizer (learning rate 0.01, momentum 0.9, weight decay 1e-3) are employed.
Task heterogeneity is more complex, as clients may work on entirely different tasks (e.g., image classification and object detection), each requiring distinct final layers, loss functions, and optimizers. Dividing the model into a shared backbone and task-specific layer can facilitate aggregation in this scenario. The CelebA dataset Liu et al. (2015) is used to evaluate task heterogeneity. It contains over 200,000 face images with up to 40 binary attributes and five landmark annotations, making it suitable for multiple tasks, including face attribute recognition and landmark detection Woubie et al. (2024); Rasouli et al. (2020); Mortaheb et al. (2022).
Two distinct task groups are defined: one performs multi-label face attribute classification, and the other conducts face landmark detection. A subset of 40% of the dataset is split evenly between the groups (in total 80%), with each group’s clients receiving equal shares, and the remaining 20% subsets of data are used for testing. Clients then partition their data into 80% training and 20% validation while sharing a common test set per group. Random rotation and horizontal flips help combat overfitting.
Both tasks employ ResNet-18 backbone and distinct task layers:
  1. 1.
    Face attribute classification uses Binary Cross Entropy loss and an AdamW optimizer 5e-5 learning rate, 5e-3 weight decay.
  2. 2.
    Face landmark detection uses Mean Squared Error loss and an Adam optimizer 5e-4 learning rate.
The experiments report four evaluation metrics: loss, precision, recall, and F1-Score. Since face landmark detection is a regression task, only the loss metric is used to evaluate this subtask. To handle class imbalance and multi-class scenarios, micro averaging strategy is used, aggregating metrics across all classes within each task group. Metrics are kept separate for different groups because their labels (in the label heterogeneity scenario) or tasks (in the task heterogeneity scenario) are distinct. All models were developed and deployed in Python using the PyTorch framework, with Lightning as the trainer and GPUs for acceleration.
4.2 Federation Setup
@@
    FedPre (intra- and cross-aggregation): Nodes use FedPre Arivazhagan et al. (2019) for both intra-group and cross-group backbone aggregation.
@@
4.3 Impact of Aggregation Frequency
@@
Table 2: Average Client Test Metrics of ColNet for Two Evaluation Scenarios with Different Local Epoch Setups
@@
Label Heterogeneity (CIFAR-10) Animal Classification 2 0.673 0.779 0.768 0.769
@@
Task Heterogeneity (CelebA) Attribute Classification 2 0.231 0.741 0.545 0.605
@@
Table 2 presents the results for four subtasks: animal classification, object classification, face attribute classification, and face landmark detection. Since face landmark detection is a regression task, only the loss metric is reported for that subtask. Besides, Figure 2 presents the epoch-wise validation loss for each subtask with different local training epoch settings.
@@
4.4 Impact of the Privatization Degree
@@
Table 3: Average Client Test Metrics of ColNet for Two Evaluation Scenarios with Different Privatization Setups
@@
Label Heterogeneity (CIFAR-10) Animal Classification BL-0 0.706 0.786 0.759 0.758
@@
Task Heterogeneity (CelebA) Attribute Classification BL-0 0.230 0.738 0.541 0.599
@@
Table 3 compares the performance of ColNet client models under these three settings in both scenarios. Besides, Figure 3 presents the validation loss in each epoch during the training process. Overall, BL-2 shows the weakest performance, implying that retaining more residual layers in the backbone benefits DFMTL by providing sufficient neurons to exchange knowledge across different tasks. At the same time, a suitable number of task layers enables proper personalization under varied conditions. Drawing on the results from both scenarios, a moderate privatization strategy—treating the last residual block and the output layer of ResNet-18 as task layers (i.e., BL-1) proves optimal and is therefore adopted for subsequent experiments.
4.5 Comparison Among Aggregation Schemes
@@
This experiment compares the performance of ColNet with (i) No aggregation (NO AGG.), (ii) Intra-aggregation (INTRA-AGG.), and (iii) FedPre with intra- and cross-aggregation (FEDPRE I&C), in label heterogeneity and task heterogeneity scenarios. Table 4 presents the test metrics of different aggregation schemes in these two scenarios after 15 federated rounds with a total of 30 epochs.
Table 4: Average Client Test Metrics for Different Aggregation Schemes
@@
Label Heterogeneity (CIFAR-10) Animal Classification NO AGG. 0.938 0.699 0.675 0.669
@@
Task Heterogeneity (CelebA) Attribute Classification NO AGG. 0.244 0.722 0.511 0.570
@@
Table 4 compares ColNet with three reference aggregation schemes under both the label heterogeneity and task heterogeneity scenarios. The results demonstrate that ColNet consistently outperforms the other methods across multiple datasets in multi-task settings.
@@
5 Conclusion and Future Work
This work introduces ColNet, a decentralized federated multi-task learning framework designed to handle both class label and task heterogeneity. ColNet organizes clients into task-specific groups, facilitating intra-group federated averaging of backbones layers. Leaders of each task group exchange the averaged backbones with leaders of other groups, employing an HCA aggregation scheme to mitigate conflicting gradients and promote effective cross-task collaboration.
Experiments across diverse datasets demonstrate its consistent superiority over other aggregation schemes. Besides, results indicate that higher aggregation frequency and a moderate degree of layer privatization could achieve performance gains.
Currently, ColNet is implemented on a single machine. Future research may investigate a fully decentralized deployment, including strategies for managing stragglers, faults, and malicious nodes, as well as dynamic clustering methods to replace predefined task groups. Additional comparisons with advanced algorithms under a truly decentralized setting could further refine ColNet’s capabilities.

### v2

@@ III-C Learning Process
[…]
8:return $B_{i}$
Model Redistribution
Group leaders broadcast the updated backbone to their members, who synchronize $B$ and continue with the next round, while task heads remain private. The rotating-leader policy is then applied to select successors for the following cross-task aggregation.
Stages (2)–(3) adapt collaboration to both model- and data-driven signals, which improves stability under label and task heterogeneity. Stage (4) integrates complementary cross-task information while mitigating gradient conflicts. The five-stage loop repeats until a stopping criterion is met.
IV Deployment and Evaluation
This section details the deployment and experimental evaluation of ColNet, covering the targeted multi-task learning scenarios, the chosen evaluation metrics, and the resulting performance outcomes  [^1] Code available at: https://github.com/Cyber-Tracer/asfdfmtl.
IV-A Scenarios, Datasets, and Models
@@
Label heterogeneity arises when clients train on different subsets of class labels, which requires splitting the model into a shared backbone and task-specific layers to protect privacy and maintain performance. Clients with the same labels form a single task group, sharing only the backbone parameters. The CIFAR-10 dataset [24] was selected to evaluate label heterogeneity, where each image belongs to one of ten classes. To simulate label heterogeneity, the dataset is partitioned into two super-classes: animals (bird, cat, deer, dog, frog, horse) and objects (airplane, automobile, ship, truck). Within each super-class, samples are split evenly and assigned to three clients in an IID manner.
@@
All clients train a ResNet-18 [25] backbone and separate task layers to accommodate different output dimensions (four classes for objects, six for animals). A Cross-Entropy loss and SGD optimizer (learning rate 0.01, momentum 0.9, weight decay 1e-3) are employed.
Task heterogeneity is more complex, as clients may work on entirely different tasks (e.g., image classification and object detection), each requiring distinct final layers, loss functions, and optimizers. Dividing the model into a shared backbone and task-specific layer can facilitate aggregation in this scenario. The CelebA dataset [26] is used to evaluate task heterogeneity. It contains over 200,000 face images with up to 40 binary attributes and five landmark annotations, making it suitable for multiple tasks, including face attribute recognition and landmark detection [27, 28, 29].
To simulate task heterogeneity, two distinct task groups are defined: multi-label face attribute classification and face landmark detection. A subset comprising 80% of the dataset is used for training/validation, created by first selecting 40% for each task group and assigning it evenly across that group’s clients; the remaining 20% serves as a common test set. Each client then partitions its local portion into 80% training and 20% validation. Standard augmentations (random rotation and horizontal flips) are applied to mitigate overfitting.
Both tasks employ ResNet-18 backbone and distinct task layers: (a) Face attribute classification uses Binary Cross Entropy loss and an AdamW optimizer 5e-5 learning rate, 5e-3 weight decay; (b) Face landmark detection uses Mean Squared Error loss and an Adam optimizer 5e-4 learning rate.
In the experiments, the overlay network uses a fully connected topology, and nodes have no knowledge of other nodes’ task types. After each local training round, each node disseminates only its backbone parameters to all other nodes in the network. Adaptive task clustering is then performed based on the received backbones and local data. Normalized spectral clustering is used with the number of clusters set to kk=2, after which intra-group aggregation and subsequent steps proceed.
The experiments report four evaluation metrics: loss, precision, recall, and F1-Score. Since face landmark detection is a regression task, only the loss metric is used to evaluate this subtask. To handle class imbalance and multi-class scenarios, micro averaging strategy is used, aggregating metrics across all classes within each task group. Metrics are kept separate for different groups because their labels (in the label heterogeneity scenario) or tasks (in the task heterogeneity scenario) are distinct.
IV-B Federation Setup
@@
    FedPre (intra- and cross-aggregation): Nodes use FedPre [16] for both intra-group and cross-group backbone aggregation.
@@
IV-C Impact of Aggregation Frequency
@@
TABLE II: Average Client Test Metrics of ColNet for Two Evaluation Scenarios with Different Local Epoch Setups
@@
Label
Heterogeneity
(CIFAR-10)
Animal
Classification
2 0.673 0.779 0.768 0.769
@@
Task
Heterogeneity
Attribute
Classification
2 0.231 0.741 0.545 0.605
@@
TABLE II presents the results for four subtasks: animal classification, object classification, face attribute classification, and face landmark detection. Since face landmark detection is a regression task, only the loss metric is reported for that subtask. Besides, Figure 2 presents the epoch-wise validation loss for each subtask with different local training epoch settings.
@@
IV-D Impact of the Privatization Degree
@@
TABLE III: Average Client Test Metrics of ColNet for Two Evaluation Scenarios with Different Privatization Setups
@@
Label
Heterogeneity
(CIFAR-10)
Animal
Classification
BL-0 0.706 0.786 0.759 0.758
@@
Task
Heterogeneity
Attribute
Classification
BL-0 0.230 0.738 0.541 0.599
@@
TABLE III compares the performance of ColNet client models under these three settings in both scenarios. Besides, Figure 3 presents the validation loss in each epoch during the training process. Overall, BL-2 shows the weakest performance, implying that retaining more residual layers in the backbone benefits DFMTL by providing sufficient neurons to exchange knowledge across different tasks. At the same time, a suitable number of task layers enables proper personalization under varied conditions. Drawing on the results from both scenarios, a moderate privatization strategy—treating the last residual block and the output layer of ResNet-18 as task layers (i.e., BL-1) proves optimal and is therefore adopted for subsequent experiments.
IV-E Comparison Among Aggregation Schemes
@@
This experiment compares the performance of ColNet with (i) No aggregation (NO AGG.), (ii) Intra-aggregation (INTRA-AGG.), and (iii) FedPre with intra- and cross-aggregation (FEDPRE I&C), in label heterogeneity and task heterogeneity scenarios. TABLE IV presents the test metrics of different aggregation schemes in these two scenarios after 15 federated rounds with a total of 30 epochs.
TABLE IV: Average Client Test Metrics for Different Aggregation Schemes
@@
Label
Heterogeneity
(CIFAR-10)
Animal
Classification
NO AGG. 0.938 0.699 0.675 0.669
@@
Task
Heterogeneity
Attribute
Classification
NO AGG. 0.244 0.722 0.511 0.570
@@
TABLE IV compares ColNet with three reference aggregation schemes under both the label heterogeneity and task heterogeneity scenarios. The results demonstrate that ColNet consistently outperforms the other methods across multiple datasets in multi-task settings.
@@
IV-F Robustness Analysis of ColNet
In the preceding experiments, all participating clients were assumed benign (honest and non-adversarial). In decentralized learning, this assumption often breaks down: some clients may behave maliciously, most notably via poisoning attacks that corrupt model updates or local data, degrade performance, and undermine model usability [30]. Accordingly, this study evaluates ColNet’s resilience to multiple poisoning settings, covering both model-poisoning and data-poisoning variants, and analyzes its model robustness.
IV-F1 Poisoning Attack Strategies
This implements four poisoning attack strategies to evaluate the model robustness of the proposed ColNet, covering data poisoning attacks (untargeted label flipping), model poisoning attacks (scaled boost and AT2FL [31]), and aggregation manipulation.
Untargeted Label Flipping. Before local training, labels are corrupted to inject label noise: for multi-label tasks, each label is flipped with probability $p$ (Bernoulli mask), and for single-label tasks, the true class is replaced with a uniformly sampled different class with probability $p$. The operation degrades overall generalization without steering predictions toward a specific alternative.
Scaled Boost. After local training but before upload, the client scales its update by a factor $s>1$ and uploads $s\cdot\Delta_{i}$, optionally preserving non-trainable statistics and applying clipping or BN re-estimation to reduce detectability, thereby amplifying the client’s influence on global updates and inducing overshoot or instability.
AT2FL (Inner-Loop Adversarial Data Poisoning). During local training, for a small number of mini-batches the attacker computes input gradients and crafts adversarial inputs $\mathbf{x}_{\text{adv}}=\mathrm{Proj}_{\mathcal{X}}(\mathbf{x}+\epsilon\,\mathrm{sign}(\nabla_{\mathbf{x}}L))$, performs additional optimization steps on these perturbed samples to steer local parameters toward directions that increase future loss on target nodes, then uploads the manipulated backbone.
Aggregation Manipulation (Malicious Aggregation Filter). At aggregation, the aggregator replaces or perturbs the conflict-averse aggregate $\Delta_{\text{HCA}}=\mathcal{A}_{\text{HCA}}(\{\Delta_{i}\})$ with $\tilde{\Delta}_{\text{agg}}=\mathcal{F}_{\text{mal}}(\Delta_{\text{HCA}},\{\Delta_{i}\})$, thereby compromising cross-task reconciliation or implanting group-level backdoors.
IV-F2 Attack evaluation setup
The robustness analysis experiments follow the same base training and evaluation protocol as the model-performance experiments. The only change is that, in each attack trial, a single client is designated as the attacker while all other clients remain benign. Comparative runs are executed in two modes: (a) clean baseline (no attacker present) and (b) attacked (one attacker active). Robustness is measured by the impact of the attack on the performance of benign clients; that is, if the loss and F1-Score of benign clients degrade only mildly under attack relative to the clean baseline, the system is considered more robust.
The dataset-specific attacker placement used in the experiments is as follows: (i) CIFAR-10 (label-heterogeneous setting). The first node in the Animal Classification (AN_N0) task is set as the attacker; the remaining nodes are benign.
(ii) CelebA (task-heterogeneous setting). The first node in the Attribute Classification (AC_N0) task is set as the attacker; the remaining nodes are benign.
IV-F3 Robustness Analysis Results
Untargeted label flipping.
This attack flips labels in the attacker’s local dataset before training. As shown in Figure 5, its effect is highly concentrated on the poisoned node: e.g., on CIFAR-10 the poisoned AN_N0’s loss increased from $\approx 0.65$ to $\approx 4.41$ and its F1 dropped from $0.77$ to $0.052$ (a $\sim 0.72$ absolute F1 loss). By contrast, same-group mates AN_N1/N2) experienced negligible aggregate harm (mean F1 stayed roughly the same: $\sim 0.734$ $\to$ $\sim 0.736$), and cross-group object nodes (OB_N) were essentially unaffected. A similar pattern holds on CelebA dataset: the poisoned attribute node (AC_N0) suffered a large F1 drop (0.591 $\to$ 0.139), while the other attribute nodes showed small or no degradation.
Figure 5: Impact of Untargeted Label Flipping Attack on CIFAR-10 and CelebA Datasets
Scaled Boost (model poisoning).
Scaled Boost amplifies a client’s update before upload (here $\times 5$ amplified). As shown in Figure 6, the poisoned client incurs the largest absolute degradation, for CIFAR-10 AN_N0 F1 fell from $0.77$ to $\approx 0.18$, but unlike simple label flips there is measurable collateral damage to same-group benign clients: AN_N1/N2 saw modest F1 drops (each $\approx 0.04\mbox{--}0.05$ absolute), while cross-group object nodes remained effectively stable (changes $\approx$ 0 or within noise). On CelebA, the attacker’s attribute node (AC_N0) lost $\sim\!0.27$ F1, and the other attribute nodes suffered smaller but non-negligible drops.
Figure 6: Impact of Scaled Boost Attack on CIFAR-10 and CelebA Datasets
AT2FL (inner-loop adversarial poisoning).
AT2FL crafts adversarial inputs during a few local mini-batches to bias the local update towards directions that maximize future loss on target nodes. As shown in Figure 7, for CIFAR-10, AN_N0’s F1 decreased from $0.77$ to $\approx 0.68$, while same-group mates (AN_N1/N2) in some runs even improved slightly, indicating that adversarially nudging shared representations can have non-uniform, task-dependent side effects. Cross-group object nodes remained essentially stable (very small changes). On CelebA the AT2FL perturbations caused small-to-moderate F1 decreases on the attribute nodes (e.g., AC_N0: $0.591\to 0.547$).
Figure 7: Impact of AT2FL Attack on CIFAR-10 and CelebA Datasets
Figure 8: Impact of Aggregation manipulation Attack on CIFAR-10 and CelebA Datasets
Aggregation manipulation.
Tampering with the aggregation routine (replacing or perturbing the conflict-averse solution) is the most system-wide attack. As shown in Figure8, in CIFAR-10 the malicious aggregation caused sizeable F1 drops across both same-group and cross-group nodes: AN nodes’ F1 decreased from $\sim 0.77$ to $\sim 0.55$ (same-group mean drop $\approx 0.18$), and object nodes’ F1 also fell substantially (cross-group mean drop $\approx 0.18$). On CelebA the aggregation attack produced roughly uniform declines across attribute nodes (each AC_N lost $\sim 0.06$ F1) and raised landmark losses.
Untargeted label flipping primarily collapses the attacker’s own model with minimal spillover to other clients; model-level manipulations (e.g., Scaled Boost) inflict stronger local damage and measurable collateral degradation within the attacker’s task group; and attacks that subvert the aggregation stage break task isolation and cause the most severe, system-wide declines, affecting both same-task and cross-task nodes.
Overall, ColNet demonstrates model robustness: it strongly resists isolated label noise and modest manipulations, but remains vulnerable when an attacker either (i) amplifies their aggregate influence or (ii) directly compromises the aggregation routine, suggesting defense priorities of anomaly detection for client updates and hardening aggregation integrity.
V Summary and Future Work
ColNet is a decentralized federated multi-task learning framework that groups clients by task, averages backbones within each group, and reconciles group updates via an HCA-style conflict-averse aggregator. An adaptive clustering module discovers and refines group membership over time. Empirical results show consistent gains over alternative aggregators while demonstrating improved robustness. The combination of a shared backbone, task-specific heads, HCA aggregation, and adaptive clustering substantially limits cross-task propagation of client-side attacks.
Future work focuses on fault-tolerant deployment, including leader election and rotation, straggler mitigation, and support for asynchronous rounds, together with stronger aggregation hardening through Byzantine-robust and secure aggregation to balance accuracy, privacy, and robustness. Large-scale distributed evaluations against stronger adaptive adversaries will quantify trade-offs and guide improvements toward comprehensive robustness guarantees.